A model rocket is shot straight up from the roof of a school. The height at any time t is approximated by the model H=15 + 23t - 5t^2,where H is the height in metres and t is the time in seconds.
a)how long does it take for the rocket to pass a window 10 m above the ground.
b) when does the rocket hit the ground?
c) what is the maximum height the rocket reaches above the roof of the school?

Answer by stanbon(75887)

A model rocket is shot straight up from the roof of a school. The height at any time t is approximated by the model H=15 + 23t - 5t^2,where H is the height in metres and t is the time in seconds.
a)how long does it take for the rocket to pass a window 10 m above the ground.
-5t^2+23t+15 = 10
-5t^2+23t+5 = 0
t = [-23 +- sqrt(23^2-4*-5*5)]/(-10)
t = [-23 +- sqrt(629)
t = [-23-25.07987..]/(-10)
t = -48.07987]/(-10)
t = 4.80798 seconds
----
Comment: The rocket is at 15 meters when it is launched.
It arrives at a height of 10 m after 4.80798 seconds on
its way down to the ground.
========================
b) when does the rocket hit the ground?
-5t^2+23t+15 = 0
t = [-23 +- sqrt(23^2-4*-1*15)]/(-10)
t = [-23 +- sqrt(829)]/(-10)
t = [-23 +- 28.7924..]/(-10)
t = 5.17924 seconds
-----------------------------
c) what is the maximum height the rocket reaches above the roof of the school?
Max occurs when t = -b/2a = -23/(-10) = 2.3 seconds
H(t) =15 + 23t - 5t^2
H(2.3) = 15 + 23*2.3 - 5(2.3)^2
H(2.3) = 41.45 m
That is 41.45 - 15 = 26.45 m above the root of the school.
==================
Cheers,
Stan H.